OpenStudy (anonymous):

solve the system y-10=3x, 27=6x+20

OpenStudy (harsimran_hs4):

take 1st eq and multiply it by 2 then subtract the equations to get the value of y what do you get ?

OpenStudy (caozeyuan):

well, from the equation 2, x=7/6, y-10=3*(7/6) so y=10+7/2=13.5

OpenStudy (anonymous):

Are you missing a y in the 2nd equation ?